OpenStudy (anonymous):

what value of a makes the equation true 24/40=9/a a.12 b.25 c.6 d.16

iYuko (iyuko):

Well, lets try each option

iYuko (iyuko):

\[\frac{ 24 }{ 40 }=\frac{ 9 }{ (12)}\]

iYuko (iyuko):

What is 24 div by 40?

OpenStudy (vhtran):

24/40=0.6

iYuko (iyuko):

Thank you

iYuko (iyuko):

Is 0.6 = to 6/(12)?

OpenStudy (vhtran):

No Problem, I didn't get an exact number when I divided the rest but you can round and see ^-^

iYuko (iyuko):

6/12 is 0.5

iYuko (iyuko):

So A is false

OpenStudy (vhtran):

I think it's 9 not 6, not sure :3

iYuko (iyuko):

@angelpooh01

iYuko (iyuko):

I need ya man...

iYuko (iyuko):

we know 24/40 is 0.6

iYuko (iyuko):

Lets try b

iYuko (iyuko):

9/25 is? @angelpooh01

iYuko (iyuko):

0.36

iYuko (iyuko):

B is false

iYuko (iyuko):

9/6 is? @angelpooh01

iYuko (iyuko):

1.5 @angelpooh01

iYuko (iyuko):

c is false

iYuko (iyuko):

9/16 @angelpooh01

OpenStudy (vhtran):

I got 0.5625, If rounded we get 0.6

iYuko (iyuko):

Right

iYuko (iyuko):

Are those the correct answer choices?

iYuko (iyuko):

The correct answer is 15

OpenStudy (anonymous):

so if 24/40 =0.6 the answer will be c

iYuko (iyuko):

If you could, please check to see if that is an answer choice

OpenStudy (anonymous):

c is my answer choice

iYuko (iyuko):

Yes

iYuko (iyuko):

C is incorrect

iYuko (iyuko):

the answer is 16

iYuko (iyuko):

@angelpooh01

OpenStudy (anonymous):

why will d be my answer if 24/40 =0.6 is short for 6 so my answer will be 6

OpenStudy (vhtran):

We are trying to replace "a" with a variable that would make the equation true. So if 24/40 is 0.6. 9/___ = 0.6? We replaced "a" (or the empty spot) with different inputs to try and get 0.6, if we input 16, we get 0.5625, when rounded you get 0.6. 16 is the closest answer with the given choices.

iYuko (iyuko):

The answer is 16

iYuko (iyuko):

We have already been through this.

OpenStudy (anonymous):

my ansswer choice d.15 it is not 16

iYuko (iyuko):

Do you understand?

iYuko (iyuko):

15 is also correct

iYuko (iyuko):

I told you that

OpenStudy (vhtran):

Ah, I see xD I think you meant to right 15 not 16. ^-^

OpenStudy (vhtran):

*write

OpenStudy (anonymous):

thank you for break can you help me with the next question please

OpenStudy (vhtran):

Sure

iYuko (iyuko):

Please make a new post

OpenStudy (anonymous):

3/8 times5/6

iYuko (iyuko):

And please select how helped you the most.

iYuko (iyuko):

0.3125

iYuko (iyuko):

0.3 is what you would type

OpenStudy (anonymous):

here is the answer choice a.1/6 b.5/16 c.4/7 d.9/20

OpenStudy (vhtran):

B

OpenStudy (vhtran):

You find the common denominator

OpenStudy (anonymous):

ok you won't mind helping me with a few more qusetions

iYuko (iyuko):

Please click Best response ----------------------------------------^^^^^^^^

OpenStudy (vhtran):

Sure

iYuko (iyuko):

----------------------------------------------------------------^^^^^^^^

OpenStudy (anonymous):

what is reciprocal of -13/8 a.13/8 b.-13/8 c.8/13 d.-8/13

OpenStudy (vhtran):

Do you know what a reciprocal is?

OpenStudy (anonymous):

no

OpenStudy (vhtran):

you basically flip the denominator with the numerator. http://www.mathsisfun.com/reciprocal.html

OpenStudy (anonymous):

it will be c.8/13

OpenStudy (vhtran):

Yes, But I'm not sure about the negative though.

OpenStudy (vhtran):

The negative stays with it. So the answer would be -8/13

OpenStudy (anonymous):

well my teacher said if there is one negative it might bea postive or a negative

OpenStudy (vhtran):

From what I have researched it should be -8/13, but depending on what your teacher said it could be either -8/13 or 8/13, You can try and take a gamble. Sorry >.<

OpenStudy (anonymous):

it is ok

OpenStudy (anonymous):

i have another question can you help me

OpenStudy (vhtran):

Is there anymore questions you need help with? I have to leave in about 8 minutes >.< I will try and answer as many questions as i can ^-^

OpenStudy (anonymous):

ok what is 2/3 divdie by 1/8 a.1/12 b.3/11 c.3/16 d.51/3

OpenStudy (vhtran):

I got 16/3

OpenStudy (vhtran):

I did the same thing as the other question, I found the common denominator and went on from there.

OpenStudy (anonymous):

ok i have 2 more question do you have time to answer it

OpenStudy (vhtran):

sure, give me both of them and I'll try to answer them ^-^

OpenStudy (anonymous):

what is the lcd of 1/6 and 3/8 a.48 b.24 c.6 d.8

OpenStudy (vhtran):

24

OpenStudy (vhtran):

I have to go, message me the question, I will be back on in a bit for a moment.

OpenStudy (vhtran):

or you can give me the question now xD

OpenStudy (anonymous):

what is 21/5+11/10 a.33/5 b.31/10 c.31/5 d.33/10

OpenStudy (vhtran):

I got 53/10 message me if you have any questions

OpenStudy (anonymous):

thank you